OpenStudy (anonymous):

What is the common difference of the arithmetic sequence -3,-0.6,1.8,4.2

OpenStudy (mathstudent55):

Subtract any number from the next one in the sequence. For example, what is 4.2 - 1.8 =?

OpenStudy (anonymous):

See, 3.6 - 1.8 = 1.8 5.4 - 3.6 = 1.8 This difference which is same is called COMMON DIFFERENCE.. Here, d= 1.8 In an Arithmetic Progression, you can get the second term, if you add COMMON DIFFERENCE in the previous term.. So, as the difference between two consecutive terms is same, So, It is an Arithmetic Progression with a Common Difference of 1.8....